# ScoreAdv: Score-based Targeted Generation of Natural Adversarial Examples via Diffusion Models

### v1

@@ Introduction
Image Recognition. While the above adversarial perturbation targets classification tasks, our method also extends to image recognition. Given a target image $\boldsymbol{x}_{tar}$ and a similarity model $f(\cdot,\cdot)$, we apply the same procedure to obtain $\bar{\boldsymbol{x}}_{t-1}$, and then sample $\tilde{\boldsymbol{x}}_{t-1}$ as follows:
$\hat{\boldsymbol{x}}_{0}$ $=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\boldsymbol{\bar{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t}}\cdot\tilde{\boldsymbol{\epsilon}}_{t-1}\right),$
where $\hat{\boldsymbol{x}}_{0}$ is the predicted clean image from $\bar{\boldsymbol{x}}_{t-1}$, and $\tilde{\boldsymbol{\epsilon}}_{t-1}$ is the predicted noise at step $t-1$. We measure similarity between $\boldsymbol{x}_{tar}$ and $\hat{\boldsymbol{x}}_{0}$ rather than $\bar{\boldsymbol{x}}_{t-1}$, since the latter contains significant noise. As the target model is trained on clean images, evaluating on denoised predictions better reflects its behavior. Thus, we first predict $\hat{\boldsymbol{x}}_{0}$ from $\bar{\boldsymbol{x}}_{t-1}$, then measure $f(\boldsymbol{x}_{tar}|\hat{\boldsymbol{x}}_{0})$.
ScoreCAM-based Inpainting
We introduce ScoreCAM to guide the diffusion denoising process with a reference image, generating natural adversarial examples that resemble it. Unlike gradient-based methods such as GradCAM (Selvaraju et al. 2017), which suffer from vanishing gradients and instability (Wang et al. 2020), ScoreCAM avoids gradient dependence.
Figure 3: Illustration of ScoreCAM.
Thus, for a given class $c$, ScoreCAM can be computed as:
We leverage the concept of image inpainting (Lugmayr et al. 2022) to perform a weighted fusion of $\boldsymbol{x}_{t-1}^{ref}$ with $\boldsymbol{\tilde{x}}_{t}$.
Noise Optimization
(Zhou et al. 2025) show that certain noise types yield better results. Motivated by this, we incorporate prior information about the target label into the initial noise. Specifically, after generating $\boldsymbol{x}_{0}$, we apply diffusion process and use gradient signals from the target model to guide noise optimization.
$\boldsymbol{x}_{T}=\left(\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}\right)+\bar{\sigma}_{T}^{2}s_{a}\nabla_{\boldsymbol{x}_{0}}\log p_{f}(y_{tar}|\boldsymbol{x}_{0}),$ (15)
$\boldsymbol{x}_{T}=\left(\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}\right)+\bar{\sigma}_{T}^{2}s_{a}\nabla_{\boldsymbol{x}_{0}}\log p_{f}(\boldsymbol{x}_{tar}|\boldsymbol{x}_{0}).$ (16)
ScoreAdv Algorithm
Experiments
Datasets
@@
Experiment Settings
We utilize the pre-trained Guided Diffusion model from OpenAI (Dhariwal and Nichol 2021), avoiding training from scratch. We select a diffusion step as $T=1000$, noise optimization iterations as $N=3$, adversarial guidance scale as $s_{a}=0.3$, and noise optimization scale as $s_{n}=0.8$.
We evaluate attacks on models from three architecture families. CNN-based models include ResNet (Res50) (He et al. 2016), VGG19 (Simonyan and Zisserman 2015), Inception-v3 (Szegedy et al. 2016), WideResNet50-2 (WRN50) (Zagoruyko and Komodakis 2016), and ConvNeXt (CNX) (Liu et al. 2022). Transformers models include ViT-B/16 (ViT-B) (Kolesnikov et al. 2021), Swin Transformer (Swin-B) (Liu et al. 2021), and DeiT-B (Touvron et al. 2020). MLP-based models include Mixer-B/16 (Mix-B) and Mixer-L/16 (Mixer-L) (Tolstikhin et al. 2021).
We use attack success rate (ASR) to measure attack effectiveness. To assess imperceptibility and visual quality, we report FID (Heusel et al. 2017), LPIPS (Zhang et al. 2018), PSNR, and SSIM (Horé and Ziou 2010).
Results
[…]
PGD 72.740 0.276 23.867 0.6541
DI2-FGSM 71.650 0.198 26.388 0.7599
S2I-FGSM 51.022 0.158 29.808 0.8296
cAdv 69.593 0.246 24.113 0.7059
NCF 70.857 0.355 14.994 0.6087
DiffAttack 48.639 0.169 27.450 0.6546
AdvDiffuser 45.399 0.137 30.146 0.8302
ScoreAdv 44.932 0.124 30.817 0.8319
@@ Results
Targeted Attack Performance. Table 1 reports the ASR on ImageNet. ScoreAdv achieves SOTA performance in most scenarios, outperforming methods such as PGD (Madry et al. 2018b), DI2-FGSM (Xie et al. 2019), S2I-FGSM (Long et al. 2022), cAdv (Bhattad et al. 2019), NCF (Yuan et al. 2022), AdvDiffuser (Chen et al. 2023a), DiffAttack (Chen et al. 2025), and Advdiff (Dai, Liang, and Xiao 2025).
While ScoreAdv may slightly underperform in isolated cases, these methods often yield lower-quality adversarial examples with poor black-box transferability. Additional results on recognition models are provided in Supplementary Material, and ScoreAdv performs even better in this setting.
Generation Quality. ScoreAdv consistently produces the most natural and coherent adversarial images. In Figure 6, it best preserves perceptual fidelity and captures high-frequency details, enhancing its realism. Table 2 presents the quantitative comparison of image quality. ScoreAdv achieves the lowest FID, indicating better preservation of the reference image distribution. Higher PSNR and SSIM reflect minimal, well-placed perturbations, while lower LPIPS confirms better alignment with human perception.
Robustness Against Defense Methods. To evaluate the robustness of ScoreAdv, we conducted a series of adversarial defense experiments. Specifically, AdvProp (Xie et al. 2020), R&P (Xie et al. 2018), RandS (Cohen, Rosenfeld, and Kolter 2019), Bit-Red (Xu, Evans, and Qi 2018), Adv-Inc-v3 (Kurakin et al. 2018b), and IncRes-v2ens (Kurakin et al. 2018a) apply defense mechanisms or train robust models before attack. SR (Mustafa et al. 2020), NRP (Naseer et al. 2020), and DiffPure (Nie et al. 2022) operate by purifying images post-attack to mitigate adversarial effects.
As shown in Table 3, ScoreAdv demonstrates superior robustness compared to previous methods and exhibits slightly better robustness than AdvDiff. This enhanced performance may be attributed to both methods that utilize the inherent distribution-shifting property of diffusion models. Moreover, the dynamic interplay of iterative denoising and adversarial perturbation injection within inadvertently contributes to improving the robustness of the generated images.
Parameter Analysis
@@
Noise Addition Steps $T^{*}$. $T^{*}$ denotes the number of diffusion steps where adversarial perturbations are applied, specifically during the final $T^{*}$ steps. As shown, increasing $T^{*}$ improves ASR but may degrade image quality due to accumulated noise. Our $T^{*}$ balances ASR and visual fidelity.
Noise Optimization Times $N$. A larger $N$ enables further optimization of the initial noise, improving adversarial effectiveness and ASR, similar to increasing iterations in GANs. However, excessive $N$ may distort the noise too much, degrading image quality. As shown in the figure, a small $N$ already yields high ASR, so we choose a moderate value to balance performance and visual fidelity.
Adversarial Guidance Scale $s_{a}$ And Noise Optimization Scale $s_{n}$. In the figure, increasing $s_{a}$ improves attack performance by amplifying perturbation strength and enhancing guidance toward the target class. However, overly large $s_{a}$ degrades image quality. The effect of $s_{n}$ follows a similar trend but has less impact, as it only applies when $N>1$.
Please see supplementary material for additional experiments on more SOTA methods and ensemble attacks and ablation studies on saliency maps.
Conclusion
In this paper, we propose ScoreAdv, an interpretable method for generating natural UAEs by leveraging the distribution shift in diffusion models. It uses a pretrained diffusion model with adversarial guidance injected during denoising, and integrates ScoreCAM to incorporate visual cues from a reference image. Experiments show that ScoreAdv achieves SOTA performance in both black-box and white-box attacks on classification and recognition tasks, while maintaining high image quality and strong robustness against defenses.
@@ Appendix A Theoretical Derivation
Appendix B Attack Performance on Recognition Systems
Table 4 presents the attack performance of various methods on the recognition model. We select IR152 (He et al. 2016), FaceNet (Schroff, Kalenichenko, and Philbin 2015), and MobileFace (Chen et al. 2018) as the target recognition model. Our proposed method, ScoreAdv, achieves SOTA performance, surpassing previous approaches such as DI (Xie et al. 2019), VMI (Wang and He 2021), SSA (Long et al. 2022), DFANet (Zhong and Deng 2021), SIA (Wang, Zhang, and Zhang 2023), BSR (Wang et al. 2024), and BPFA (Zhou et al. 2024). As shown in the table, ScoreAdv significantly outperforms other methods in black-box scenarios and achieves competitive or superior results in most white-box settings. This demonstrates that our method is not only effective in attacking image classification models but also exhibits strong performance when targeting recognition models.
Appendix C Additional Results
Comparison With More SOTA Methods
While we have already compared our method with the most advanced diffusion-based UAE approaches, we further extend our evaluation to include additional SOTA methods, such as TI-FGSM (Dong et al. 2019) VENOM (Kuurila-Zhang, Chen, and Zhao 2025), AdvAD (Li et al. 2024), and NatADiff (Collins et al. 2025). As shown in Table 5, ScoreAdv consistently achieves superior performance in both attack success rate and image quality. Moreover, although comparing ScoreAdv with ensemble-based attacks is inherently disadvantageous to us, we still include such comparisons. The results demonstrate that our method remains highly competitive against a wide range of baselines.
@@ Appendix C Additional Results
Ablation Study on Saliency Map
As shown in Table 6, we compare the proposed ScoreCAM with other interpretability-based saliency methods, including GradCAM (Selvaraju et al. 2017), and GradCAM++ (Chattopadhyay et al. 2018). Experimental results demonstrate that our approach consistently outperforms the others in both attack effectiveness and image quality.

### v2

@@ 4.1 Datasets
4.2 Experiment Settings
We utilize the pre-trained Guided Diffusion model from OpenAI [12], avoiding training from scratch. We select a diffusion step as $T=1000$, noise optimization iterations as $N=3$, adversarial guidance scale as $s_{a}=0.3$, and noise optimization scale as $s_{n}=0.8$.
We evaluate attacks on models from three architecture families. CNN-based models include ResNet (Res50) [17], VGG19 [43], Inception-v3 [46], WideResNet50-2 (WRN50) [58], and ConvNeXt (CNX) [29]. Transformers models include ViT-B/16 (ViT-B) [23], Swin Transformer (Swin-B) [28], and DeiT-B [48]. MLP-based models include Mixer-B/16 (Mix-B) and Mixer-L/16 (Mixer-L) [47].
We use attack success rate (ASR) to measure attack effectiveness. To assess imperceptibility and visual quality, we report FID [18], LPIPS [60], PSNR, and SSIM [21], which are all computed between the generated adversarial images and the reference images.
4.3 Results
@@
Targeted Attack Performance.
Table 1 presents the ASR ofdifferent methods on the ImageNet dataset. We compare ScoreAdv with previous advanced adversarial attack methods such as PGD [32], DI2-FGSM [54], S2I-FGSM [30], cAdv [1], NCF [57], AdvDiffuser [6], DiffAttack [3], and Advdiff [11]. Apart from potential resolution adjustments, all implementations remain consistent with their settings.
As shown in Table 1, ScoreAdv achieves SOTA performance in nearly all black-box and white-box scenarios. While ScoreAdv may slightly underperform in certain cases, these methods often yield lower-quality adversarial examples with poor black-box transferability. Additional results on recognition models are provided in Supplementary Material 7. Both quantitative and qualitative results show that ScoreAdv performs even better in this setting.
Generation Quality.
[…]
PGD 72.740 0.276 23.867 0.6541
DI2-FGSM 71.650 0.198 26.388 0.7599
S2I-FGSM 51.022 0.158 29.808 0.8296
cAdv 69.593 0.246 24.113 0.7059
NCF 70.857 0.355 14.994 0.6087
DiffAttack 48.639 0.169 27.450 0.6546
AdvDiffuser 45.399 0.137 30.146 0.8302
ScoreAdv 44.932 0.124 30.817 0.8319
ScoreAdv consistently produces the most natural and coherent adversarial images. In Figure 6, it best preserves perceptual fidelity and captures high-frequency details, enhancing its realism. Table 3 presents the quantitative comparison of image quality. ScoreAdv achieves the lowest FID, indicating better preservation of the reference image distribution. Higher PSNR and SSIM reflect minimal, well-placed perturbations, while lower LPIPS confirms better alignment with human perception.
Robustness Against Defense Methods.
To evaluate the robustness of ScoreAdv, we conducted a series of adversarial defense experiments. Specifically, AdvProp [55], R&P [53], RandS [9], Bit-Red [56], Adv-Inc-v3 [25], and IncRes-v2ens [24] apply defense mechanisms or train robust models before attack. SR [35], NRP [36], and DiffPure [38] operate by purifying images post-attack to mitigate adversarial effects.
As shown in Table 2, ScoreAdv demonstrates superior robustness compared to previous methods and exhibits slightly better robustness than AdvDiff. This enhanced performance may be attributed to both methods that utilize the inherent distribution-shifting property of diffusion models. Moreover, the dynamic interplay of iterative denoising and adversarial perturbation injection inadvertently contributes to improving the robustness of the generated images.
Computational Analysis
[…]
Additional Results.
Please see Supplementary Material 8 for additional experiments on more SOTA methods and ensemble attacks and ablation studies on saliency maps.
4.4 Parameter Analysis
@@
Noise Addition Steps $T^{*}$.
$T^{*}$ denotes the number of diffusion steps where adversarial perturbations are applied, specifically during the final $T^{*}$ steps. As shown, increasing $T^{*}$ improves ASR but may degrade image quality due to accumulated noise. Our $T^{*}$ balances ASR and visual fidelity.
Noise Optimization Times $N$.
A larger $N$ enables further optimization of the initial noise, improving adversarial effectiveness and ASR, similar to increasing iterations in GANs. However, excessive $N$ may distort the noise too much, degrading image quality. As shown in the figure, a small $N$ already yields high ASR, so we choose a moderate value to balance performance and visual fidelity.
Adversarial Guidance Scale $s_{a}$ And Noise Optimization Scale $s_{n}$.
In the figure, increasing $s_{a}$ improves attack performance by amplifying perturbation strength and enhancing guidance toward the target class. However, overly large $s_{a}$ degrades image quality. The effect of $s_{n}$ follows a similar trend but has less impact, as it only applies when $N>1$.
5 Conclusion
In this paper, we propose ScoreAdv, an interpretable method for generating natural UAEs by leveraging the distribution shift in diffusion models. It uses a pretrained diffusion model with adversarial guidance injected during denoising, and integrates ScoreCAM to incorporate visual cues from a reference image. Experiments show that ScoreAdv achieves SOTA performance in both black-box and white-box attacks on classification and recognition tasks. Crucially, it manages to maintain exceptionally high image quality and exhibits strong robustness against defensive mechanisms, all while remaining computationally efficient. Moreover, since ScoreAdv only relies on the gradient of a generic scoring function, its formulation is naturally extensible to vision-language models by substituting $f$ with multimodal encoders. This flexibility further highlights its generality beyond conventional image classifiers.
@@ 6.3 Derivation of equation (15)
7 Attack Performance on Recognition Systems
Table 4 presents the attack performance of various methods on the recognition model. We select IR152 [17], FaceNet [41], and MobileFace [5] as the target recognition model. Our proposed method, ScoreAdv, achieves SOTA performance, surpassing previous approaches such as DI [54], VMI [51], SSA [30], DFANet [65], SIA [52], BSR [50], and BPFA [66]. As shown in the table, ScoreAdv significantly outperforms other methods in black-box scenarios and achieves competitive or superior results in most white-box settings. This demonstrates that our method is not only effective in attacking image classification models but also exhibits strong performance when targeting recognition models.
8 Additional Results
8.1 Comparison With More SOTA Methods
While we have already compared our method with the most advanced diffusion-based UAE approaches, we further extend our evaluation to include additional SOTA methods, such as TI-FGSM [14] VENOM [26], AdvAD [27], and NatADiff [10]. As shown in Table 5, ScoreAdv consistently achieves superior performance in both attack success rate and image quality. Moreover, although comparing ScoreAdv with ensemble-based attacks is inherently disadvantageous to us, we still include such comparisons. The results demonstrate that our method remains highly competitive against a wide range of baselines.
@@ 8.1 Comparison With More SOTA Methods
8.2 Ablation Study on Saliency Map
As shown in Table 6, we compare the proposed ScoreCAM with other interpretability-based saliency methods, including GradCAM [42], and GradCAM++ [2]. Experimental results demonstrate that our approach consistently outperforms the others in both attack effectiveness and image quality.